### v4

# REArtGS: Reconstructing and Generating Articulated Objects via 3D Gaussian Splatting with Geometric and Motion Constraints

Abstract
Articulated objects, as prevalent entities in human life, their 3D representations play crucial roles across various applications. However, achieving both high-fidelity textured surface reconstruction and dynamic generation for articulated objects remains challenging for existing methods. In this paper, we present REArtGS, a novel framework that introduces additional geometric and motion constraints to 3D Gaussian primitives, enabling realistic surface reconstruction and generation for articulated objects. Specifically, given multi-view RGB images of arbitrary two states of articulated objects, we first introduce an unbiased Signed Distance Field (SDF) guidance to regularize Gaussian opacity fields, enhancing geometry constraints and improving surface reconstruction quality. Then we establish deformable fields for 3D Gaussians constrained by the kinematic structures of articulated objects, achieving unsupervised generation of surface meshes in unseen states. Extensive experiments on both synthetic and real datasets demonstrate our approach achieves high-quality textured surface reconstruction for given states, and enables high-fidelity surface generation for unseen states. Project site: https://sites.google.com/view/reartgs/home.
1 Introduction
Articulated objects are ubiquitous in our daily lives. Modeling articulated objects, i.e. mesh reconstruction and generation, holds significant importance in many fields of computer vision and robotics fields including virtual and augmented reality liu2022toward ; xue2021omad , object manipulation yu2024gamma ; wang2024rpmart and human-object interaction yang2022oakink ; li2024semgrasp . Currently, the surface and shape reconstruction for articulated objects is a non-trivial task due to the challenges: firstly, articulated objects exhibit complicated and diverse geometric structures with a wide range of scales. Secondly, articulated objects possess varied kinematic structures, and the static surface meshes generated from vanilla 3D reconstruction methods fail to meet practical interaction requirements. Under this circumstance, PARIS liu2023paris attempts to use multi-view images from two states for articulated object dynamically reconstruction with neural implicit radiance fields. Nevertheless, PARIS lacks geometric constraints, leading to shape-radiance ambiguity and additional errors for motion analysis.
Figure 1: Given multi-view RGB images of articulated objects from two arbitrary states, our REArtGS enables high-quality textured surface mesh reconstruction and generation for unseen states.
In recent years, 3D Gaussian Splatting (3DGS) kerbl20233d achieves realistic and real-time novel view synthesis through explicit representation with 3D Gaussian primitives. Some subsequent works have expanded 3DGS to surface reconstruction yu2024gaussian ; huang20242d ; chen2024pgsr and dynamic reconstruction  yang2024deformable ; liu2024dynamic ; duan20244d . However, existing 3DGS-based surface reconstruction methods typically suffer from insufficient geometric constraints yu2024gaussian or impose constraints by restricting the shapes of Gaussian primitives  huang20242d ; guedon2024sugar , leading to noisy surface reconstruction results. On the other hand, current dynamic reconstruction methods via 3DGS tend to input motion time and spatial positions into neural networks to obtain the deformed positions of Gaussian primitives yang2024deformable ; duan20244d . Consequently, these methods require continuous supervision throughout the entire motion, which limits their ability to generate surface meshes in unseen states. In general, introducing 3DGS as a ready-to-use technique into articulated object surface reconstruction and generation is still not feasible and remains a challenging task.
In this paper, to address the aforementioned issues and make full advantage of 3DGS, we propose REArtGS, a novel approach that Reconstructs and gEnerates high-quality texture-rich mesh surfaces for Articulated objects via 3D Gaussian Splatting, only taking multi-view RGB images from two arbitrary states. Specifically, unlike most 3DGS methods based on the opacity field, which may lead to noisy surface extraction due to the non-strict linearity of the opacity field, we first introduce the Signed Distance Function (SDF) field to facilitate geometry learning. Then we propose a geometry constraint which encourages the SDF values of the Gaussian primitives approach zero when their opacity values reach the maximum , achieving unbiased guidance of the SDF field over the opacity field. In this way, we can take advantage of the strong linearity of the SDF field to explicitly establish a connection between Gaussian opacity field and the scene surface, which significantly enhances geometric learning and reduces artifacts.
Subsequently, we employ the optimized Gaussian primitives as an accurate geometry initialization for dynamic surface generation. Due to the absence of intermediate states, we leverage the kinematic structures of articulated objects to model time-continuous Gaussian deformable fields in an unsupervised fashion, and constrain the deformable fields using learnable motion parameters. Concretely, we propose a heuristic method for unsupervised part segmentation and formulate the deformable fields of dynamic parts via the motion parameters. Our REArtGS is evaluated on PartNet-Mobility xiang2020sapien and AKB-48 liu2022akb object repositories, ranging from synthetic to real-world data. Extensive experiments demonstrate that our REArtGS outperforms existing state-of-the-art methods in articulated object surface reconstruction and generation tasks, as shown in Fig. 1.
In summary, our main contributions can be summarized as follows: (1) We propose REArtGS, a novel framework introducing 3DGS to conduct high-quality textured surface reconstruction and time-continuous generation for articulated objects, only using multi-view images from two arbitrary states. (2) Our REArtGS exploits an unbiased SDF guidance for 3D Gaussian primitives to enhance geometric constraints for improving reconstruction quality, and also establishes the deformable fields constrained by kinematic structures of articulated objects to generate unseen states in an unsupervised manner. (3) We incorporate REArtGS into various scenes ranging from synthetic to the real-world data across many different articulation categories. The extensive experimental results demonstrate that our approach significantly outperforms SOTAs in both mesh reconstruction and generation tasks.
2 Related Work
2.1 Articulated Object Shape Reconstruction
Object surface reconstruction is a well-established problem for understanding the full geometric shape of objects. Some works introduce to encode continuous functions that model the objects using Signed Distance park2019deepsdf ; mu2021sdf , radiance mildenhall2021nerf ; rakotosaona2024nerfmeshing ; tang2023delicate and occupancy mescheder2019occupancy . To achieve both geometry and motion analysis in a single forward, Ditto jiang2022ditto and REACTO song2024reacto propose to generate shapes at unseen states from pair observations. Specifically, PARIS liu2023paris alleviates the 3D data requirement limitation and succeeds in reconstructing shape surface with only RGB images. Most recently, ArticulatedGS guo2025articulatedgs and ArtGS liu2025building introduce 3DGS to achieve both the reconstruction and motion estimation of articulated objects. However, these methods still suffer from insufficient geometry constraints, which may lead to noisy surface mesh outputs. Therefore, our REArtGS aims to address these issues, and proposes geometric and motion constrained 3DGS for this task.
2.2 Surface Reconstruction with 3DGS
3D Gaussain Splatting has become increasingly popular technique for surface reconstruction in recent years kerbl20233d ; li2024loopgaussian ; yu2024mip ; qin2024langsplat . GOF yu2024gaussian establishes opacity fields of 3D Gaussians using ray-tracing-based rendering, and extracts the surface meshes by the opacity level set. 2DGS huang20242d and PGSR chen2024pgsr seek to transform 3D Gaussians into 2D flat representation, obtaining accurate normal distribution. Although improving the surface reconstruction quality, these methods still lack more reasonable geometry constraints. Several researchers attempt to integrate SDF representation with 3D Gaussians yu2024gsdf ; chen2023neusg ; xiang2024gaussianroom ; dai2024high , but these approaches commonly use SDF to regularize normals and guide the pruning of 3D Gaussians, without substantial optimization of Gaussian opacity fields. Meanwhile, some works conduct dynamic reconstruction using 3DGS. Deformable 3DGS proposes a deformation field to reconstruct dynamic scenes with 3D Gaussian primitives. 4DGS duan20244d leverages a spatial-temporal structure encoder and a multi-head Gaussian deformation decoder to derive the deformed 3D Gaussians at a given time. However, these methods rely on complete supervision of the motion process, limiting their capacity for generating unseen states. The similar issues also occur in DGMesh liu2024dynamic and REACTO song2024reacto . We provide more analysis and comparison of the related works in Appendix.
Figure 2: The overall pipeline of REArtGS. We introduce additional geometric and motion constraints for 3D Gaussian primitives, achieving high-quality surface mesh reconstruction and time-continuous generation, with only multi-view images from arbitrary two states.
3 Method
The overall framework of our REArtGS is illustrated in Fig. 2. Taking multi-view RGB images from two arbitrary states $s=0$, $s=1$ of articulated objects, we aim to achieve high-quality textured mesh reconstruction and generation at any unseen states between $s\in[0,1]$.
We first introduce SDF representation and propose an unbiased SDF regularization to enhance the geometry constraints of 3D Gaussian primitives. In this manner, we improve the reconstruction quality and yield dense point clouds at state $s=0$, providing an accurate geometric prior for subsequent dynamic reconstruction. Then we establish time-continuous deformable fields for 3D Gaussian primitives constrained by the kinematic structures of articulated objects. Given any state $s$, we can derive the deformed position $x_{s}$ of Gaussian primitives through the deformation fields in an unsupervised fashion. The details of our approach are elaborated below.
Figure 3: The illustration of unbiased SDF regularization. It can be observed that when $t$ approaches $t^{*}$, the absolute value of SDF $|\mathcal{S}|$ converges to zero with the unbiased SDF regularization.
3.1 Reconstruction with Unbiased SDF Guidance
To better preserve the geometric features of Gaussian primitives, we use the ray-tracing-based 3DGS following GOF, which evaluates rendering contribution of Gaussian primitives directly without the 3D-to-2D projection step. However, the Gaussian primitives still suffer from insufficient geometric constraints. Concretely, the opacity fields lack an explicit link with scene surface, which may lead to noisy outputs in surface reconstruction of articulated objects.
To address the limitation, we first introduce SDF representation to guide the geometry learning of Gaussian primitives. We utilize a Multi-layer Perceptrons (MLP) with 8 hidden layers to learn the SDF values for spatial position inputs, and the scene surface can be represented by the zero-set:
$\mathcal{S}=\left\{\mathbf{x}\in\mathbb{R}^{3}\mid f(\mathbf{x})=0\right\}$ (1)
Following Neus wang2021neus , we can derive the opacity $\bm{\hat{\sigma}}$ for the center $\mathbf{x}_{i}$ of Gaussian $G_{i}$ from learned SDF values:
$\bm{\hat{\sigma}}_{i}=\text{max}\left(\frac{\Phi\left(f\left(\mathbf{x}_{i}\right)\right)-\Phi\left(f\left(\mathbf{x}_{i+1}\right)\right)}{\Phi\left(f\left(\mathbf{x}_{i}\right)\right)},0\right)$ (2)
where $\Phi$ denotes a Sigmoid function. Nevertheless, the SDF-opacity conversion is not directly applicable to the $\alpha$ blending rendering. This is because each 3D Gaussian primitive models a local density distribution, while Eq. 2 only maps the global positions to their opacity values and neglects the local properties of Gaussian primitives. Therefore, we define the opacity of Gaussian primitive combining the rendering contribution $\bm{\varepsilon}$ proposed in GOF as:
$\bm{\sigma_{i}}$ $=\bm{\hat{\sigma}}_{i}\bm{\varepsilon}\left(G_{i}\right)$ (3)
$=\bm{\hat{\sigma}}_{i}\cdot\text{max}(e^{-\frac{1}{2}{\mathbf{x}_{L}}^{T}\mathbf{x}_{L}})$
$=\bm{\hat{\sigma}}_{i}\cdot\text{max}\left(e^{-\frac{1}{2}\left(\mathbf{r}_{L}^{T}\mathbf{r}_{L}t^{2}+2\mathbf{o}_{L}^{T}\mathbf{r}_{L}t+\mathbf{o}_{L}^{T}\mathbf{o}_{L}\right)}\right)$
where $\mathbf{o}_{L}$, $\mathbf{r}_{L}$ are the camera center and incident ray direction represented in the local coordinate system of Gaussian $G_{i}$, which can be acquired by local scaling matrix and rotation matrix of $G_{i}$. The maximum value of $\bm{\varepsilon}\left(G_{i}\right)$ occurs at ray depth $t^{*}$, expressed as follows:
$t^{*}=\frac{\mathbf{o}_{L}^{T}\mathbf{r}_{L}}{\mathbf{r}_{L}^{T}\mathbf{r}_{L}}$ (4)
The rendering equation based on $\bm{\sigma}$ is represented as:
$\mathbf{C}=\sum_{i=1}^{N}\mathbf{c}_{i}\bm{\sigma_{i}}\prod_{j=1}^{i-1}\left(1-\bm{\sigma_{i}}\right)$ (5)
where $N$ is the number of Gaussian primitives involved in $\alpha$-blending and $\mathbf{c}_{i}$ is the color modeled with spherical harmonics. To leverage SDF to constrain the geometric learning of Gaussian primitives, we first introduce a bell-shaped function $\Phi_{k}$ to modulate the transformation from SDF to opacity and replace the original Sigmoid activation function $\Phi$ in Eq. 2, formulated as:
$\Phi_{k}(f(\mathbf{x}))=\frac{e^{k\cdot f(\mathbf{x})}}{\left(1+e^{k\cdot f(\mathbf{x})}\right)^{2}}$ (6)
where $k$ is a learnable parameter that adjusts the function shape and we initially set it to 0.1. Intuitively, the Gaussian primitive closer to the surface has a higher opacity value.
However, since $\bm{\hat{\sigma}}$ is determined solely by the Gaussian’s center, a non-alignment still exists between the maximum points of $\bm{\hat{\sigma}}$ and $\bm{\varepsilon}$, which can be formulated as:
$t_{\text{bias}}=\left\|\underset{t}{\text{argmax}}(\bm{\hat{\sigma}})-\underset{t}{\text{argmax}}(\bm{\varepsilon)}\right\|$ (7)
To mitigate this bias, we propose an unbiased regularization for the SDF value at depth $t^{*}$ as following:
$\mathcal{L}_{\text{unbias}}=\left\|f(\mathbf{o}+t^{*}\mathbf{r})\right\|_{2}^{2}$ (8)
where $\mathbf{o}$, $\mathbf{r}$ are the camera center and incident ray direction. Through the unbiased regularization, we encourage that when the rendering contribution of a Gaussian primitive reaches its maximum value, the corresponding spatial position is close to the scene surface, as shown in Fig. 3. In this manner, the SDF representation is able to regularize the opacity fields without bias and facilitate more reasonable distribution of Gaussian primitives over scene surface. We provide more detailed elaboration and proof in the Appendix.
3.2 Mesh Generation with Motion Constraints
Due to the explicit representation of Gaussian primitives, we can acquire accurate dense point clouds from reconstruction, serving as initial geometry for mesh generation. Then we establish time-continuous deformable fields for 3D Gaussians to infer the deformed position $\mathbf{x}_{s}$ at state $s$. However, it is challenging to optimize the deformable fields for Gaussian primitives only with the supervision of two states. Accordingly, we exploit the kinematic structures of articulated objects to constrain the Gaussian deformation field. We first assume that the articulated motion only includes rotation and translation following PARIS. Then we employ a heuristic method to efficiently predict the joint type through movement trend of points, which is elaborated in the Appendix. For rotation joints, our learning objective is to determine their rotation axis and rotation angle. Specifically, we take the pivot point $\mathbf{o}_{r}\in\mathbb{R}^{3}$ of the rotation axis and a normalized quaternion $\mathbf{q}\in\mathbb{R}^{4}$ as learnable parameters, where the quaternion can be decoupled into the rotation axis $\mathbf{a}\in\mathbb{R}^{3}$ and rotation angle $\theta$. For prismatic joints, our learning objective is to determine their translation directions $\mathbf{d}$ and translation distances $m$. Similarly, we take the unit vector $\mathbf{d}\in\mathbb{R}^{3}$ representing the translation direction and the translation distance $m$ as learnable parameters.
Subsequently, we formulate the deformable fields through a canonical state. Given the rotation motion $SO(3)\in\mathbb{R}^{3\times 3}$, we define its canonical state as $s^{*}=0.5$ and the rotation angle of state $s$ can be expresses as:
$\theta_{s}=\frac{\left(s^{*}-s\right)}{s^{*}}\theta$ (9)
where $s\in\left[0,1\right]$ and $\theta\in\left[-\pi/2,\pi/2\right]$. The angle-bounded parameterization is able to prevent the singularity in exponential coordinates when $\left\|\theta\right\|>\pi$. Using Rodrigues’ rotation formula pina2011rotations , the deformed position $\mathbf{x}_{s}$ can be derived as following:
$\mathbf{x}_{s}=\left(\mathbf{I}+\sin(\theta_{s})\mathbf{K}+(1-\cos(\theta_{s}))\mathbf{K}^{2}\right)\left(\mathbf{x}-\mathbf{o}_{r}\right)+\mathbf{o}_{r}$ (10)
where $\mathbf{K}$ is a skew-symmetric matrix formed by the rotation rotation axis $\mathbf{a}$ and $\mathbf{I}\in\mathbb{R}^{3\times 3}$ is a unit matrix. Meanwhile, we align the local quaternions of 3D Gaussian primitives after rotation through quaternion multiplication.
For the prismatic motion $m\cdot\mathbf{d}\in\mathbb{R}^{3}$, we take advantage of the vector space structure of Euclidean geometry by setting $s^{*}=0$. The deformed position $\mathbf{x}_{s}$ can be naturally acquired by linear interpolation, formulated as:
$\mathbf{x}_{s}=\mathbf{x}+s\cdot m\cdot\mathbf{d}$ (11)
unlike rotation motion, this parameterization avoids singularities due to the flat Riemannian structure of $\mathbb{R}^{3}$.
However, the deformable fields should be applied exclusively to movable Gaussian primitives. Consequently, the Gaussian primitives need to be segmented into dynamic and static parts. To address this challenge, we propose an unsupervised method for dynamic part segmentation. We first only input the multi-view images at the end state during warm-up training iterations, and then perform an initialization segmentation of the dynamic part. The segmentation criterion for the initial dynamic components is represented as: $\left|\Delta_{\mathbf{x}}\right|>\beta$, where $\left|\Delta_{\mathbf{x}}\right|$ is the L1 distance of spatial position variation during the warm-up training, and $\beta$ is its average value. Furthermore, the spatial transformation of the dynamic part should align with the learnable motion parameters. Hence the dynamic part among the Gaussian primitives is re-evaluated at certain intervals of iterations, where the criterion for revolute or prismatic joints is represented as:
$\left|\hat{\theta}-\theta\right|$ $<\frac{\varphi_{\theta}}{K}$ (12)
$\left|\hat{m}-m\right|$ $<\frac{\varphi_{m}}{K}$
where $\hat{\theta}$, $\hat{m}$ are the rotation angle around axis $\mathbf{a}$ and the translation distance respectively, $\varphi$ is the tolerance threshold which decreases as the number of iterations $K$ increases.
Moreover, our method can be conveniently transferred to multi-part articulated objects, through sequentially learning the segmentation masks and motions of each individual part. Please refer to our Appendix for more details.
3.3 Optimization and Textured Mesh Extraction
Figure 4: The qualitative result of surface reconstruction on PartNet-Mobility dataset. We show both textured and non-textured meshes for the best comparison.
Since the shapes of Gaussian primitives are not limited, the normal distributions of Gaussian primitives are difficult to estimate accurately. However, Gaussian’s normals should be parallel to the gradients of SDF in an ideal case. We consequently introduce additional regularization to regularize the normals of Gaussian primitives by SDF representation, expressed as follows:
$\mathcal{L}_{\text{normal}}=\frac{1}{N}\sum_{i}^{N}(1-\frac{\left|\mathbf{n}_{i}\cdot\nabla f(\mathbf{x}_{i})\right|}{\left\|\mathbf{n}_{i}\right\|\cdot\left\|f(\mathbf{x}_{i})\right\|})$ (13)
where $\mathbf{n}_{i}$ represents the normal of Gaussian $G_{i}$, which is calculated following GOF:
$\mathbf{n}_{i}=-R_{i}^{T}S_{i}^{-1}\mathbf{r}_{L}$ (14)
where $R$ is the rotation matrix encoded by Gaussian’s quaternion and $S$ is the scaling matrix. We also employ the Eikonal regularization term gropp2020implicit to encourage SDF gradient direction perpendicular to the surface, represented as:
$\mathcal{L}_{\text{eik}}=\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\nabla f(\mathbf{x}_{i}))\right\|_{2}-1\right)^{2}$ (15)
The overall training objective $\mathcal{L}$ is formulated as follows:
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{\text{unbias}}+\lambda_{2}\mathcal{L}_{\text{normal}}+\lambda_{3}\mathcal{L}_{\text{eik}}+\lambda_{3}\mathcal{L}_{d}$ (16)
where $\lambda$ is the weight of regularization and $\mathcal{L}_{d}$ is the depth distortion loss following huang20242d . $\mathcal{L}_{c}$ is defined as:
$\mathcal{L}_{c}=\lambda\mathcal{L}_{\text{D-SSIM}}+(1-\lambda)\mathcal{L}_{1}$ (17)
where $\mathcal{L}_{\text{D-SSIM}}$ is proposed in kerbl20233d and $\mathcal{L}_{1}$ is the L1 norm of the pixel loss.
Once the optimization converges, we adopt the TSDF fusion algorithmnewcombe2011kinectfusion to extract the textured mesh from Gaussian primitives. We utilize the $\alpha$-blending rendering pipeline to obtain depth, opacity and RGB renderings for training views. Then we integrate these images into a voxel block grid (VBG) and extract a triangle mesh from the VBG. The textures of surface meshes can be efficiently obtained leveraging the spherical harmonics and opacity of Gaussian primitives. Please refer to our Appendix for a more detailed illustration.
Table 1: Quantitative results for the surface reconstruction quality on PartNet-Mobility dataset. We bold the best results and underline the second best results. * means we implement it without depth supervision for fair comparison.
Metrics Method Stapler USB Scissor Fridge Foldchair Washer Blade Laptop Oven Storage Mean
CD(ws) $\downarrow$ A-SDF mu2021sdf 14.19 7.14 10.61 13.71 40.85 12.50 3.31 2.11 21.37 22.57 14.84
Ditto jiang2022ditto 2.38 2.09 1.70 2.16 6.80 7.29 42.04 0.31 2.51 3.91 7.19
PARIS liu2023paris 0.96 1.80 0.30 2.68 0.42 18.31 0.46 0.25 6.07 8.12 3.94
GOF yu2024gaussian 1.51 8.09 0.36 1.51 0.53 17.35 0.89 0.84 19.81 10.04 6.09
ArtGS liu2025building 2.77 1.36 0.75 2.01 0.41 20.59 0.63 0.99 9.01 9.00 4.75
REArtGS (Ours) 3.47 0.75 0.29 1.50 0.40 12.20 0.72 0.53 8.89 8.27 3.79
CD(rs) $\downarrow$ A-SDF park2019deepsdf 2.140 2.478 1.805 1.801 2.080 1.590 5.630 2.418 2.751 4.077 2.677
Ditto jiang2022ditto 2.874 3.049 2.926 1.550 0.925 1.428 5.792 1.184 1.296 2.837 2.386
PARIS liu2023paris 2.077 3.812 2.807 0.370 1.209 3.869 2.855 1.025 2.873 2.903 2.380
GOF yu2024gaussian 2.510 1.736 2.304 2.230 2.787 5.725 2.111 0.961 3.080 2.812 2.626
ArtGS liu2025building 2.949 1.704 2.438 0.725 0.615 2.874 2.031 1.182 1.148 1.157 1.682
REArtGS (Ours) 2.186 1.433 2.291 0.475 0.018 1.204 2.596 0.038 0.784 1.330 1.236
F1 $\uparrow$ A-SDF park2019deepsdf 0.041 0.035 0.094 0.019 0.053 0.046 0.224 0.001 0.010 0.015 0.054
Ditto jiang2022ditto 0.197 0.181 0.275 0.114 0.352 0.059 0.107 0.366 0.052 0.030 0.173
PARIS liu2023paris 0.240 0.151 0.343 0.091 0.429 0.024 0.396 0.533 0.031 0.033 0.227
GOF yu2024gaussian 0.217 0.189 0.614 0.173 0.480 0.087 0.386 0.303 0.049 0.036 0.253
ArtGS liu2025building 0.251 0.209 0.438 0.118 0.447 0.053 0.408 0.294 0.056 0.042 0.232
REArtGS (Ours) 0.256 0.307 0.598 0.165 0.502 0.069 0.488 0.419 0.065 0.066 0.294
EMD $\downarrow$ A-SDF park2019deepsdf 0.755 1.113 0.952 0.945 1.020 0.982 1.763 1.039 1.174 1.258 1.100
Ditto jiang2022ditto 1.724 1.308 1.212 0.619 0.935 0.841 1.996 0.417 0.852 0.970 1.087
PARIS liu2023paris 1.197 1.381 0.637 0.426 0.778 1.660 1.823 0.706 1.200 1.046 1.085
GOF yu2024gaussian 1.643 1.815 1.075 1.055 1.181 1.692 1.237 0.394 1.241 0.959 1.229
ArtGS liu2025building 1.921 0.677 0.934 0.637 0.506 1.154 1.131 0.502 1.006 0.787 0.926
REArtGS (Ours) 1.060 0.847 1.078 0.485 0.097 0.777 1.112 0.111 0.627 0.755 0.695
Figure 5: The qualitative results of surface generation at arbitrary unseen states on PartNet-Mobility dataset. We show both textured and non-textured meshes for best comparison. The states are sampled randomly.
Table 2: Quantitative results for the surface generation quality on PartNet-Mobility dataset. We bold the best results and underline the second best results. * means we implement it without depth supervision for fair comparison.
Metrics Method Stapler USB Scissor Fridge Foldchair Washer Blade Laptop Oven Storage Mean
CD(ws) $\downarrow$ A-SDF mu2021sdf 112.30 34.01 94.11 33.46 56.73 55.82 3.78 27.64 29.60 25.30 47.28
PARIS liu2023paris 2.09 14.51 12.81 2.64 10.42 18.30 0.80 3.25 10.89 5.89 8.16
D-3DGS yang2024deformable 98.90 16.79 91.84 48.78 30.10 30.93 1.262 65.11 22.39 18.79 42.49
ArtGS liu2025building 2.34 1.72 0.68 2.05 0.52 20.59 0.63 1.00 8.78 10.97 4.92
REArtGS (Ours) 2.60 1.68 0.44 2.21 0.41 14.80 0.59 3.65 8.11 6.48 4.10
CD(rs) $\downarrow$ A-SDF mu2021sdf 8.992 1.806 1.822 5.166 5.668 7.050 5.991 2.710 5.245 7.794 5.224
PARIS liu2023paris 2.113 0.417 4.704 0.408 1.216 0.857 4.442 1.354 2.793 4.332 2.264
D-3DGS yang2024deformable 9.589 1.544 0.891 0.465 1.147 3.218 0.168 0.449 3.773 3.100 2.434
ArtGS liu2025building 4.924 0.670 0.381 0.471 1.622 2.670 1.455 1.269 3.912 0.562 1.794
REArtGS (Ours) 3.651 0.487 0.127 0.464 0.420 1.617 0.414 0.298 2.129 1.020 1.063
F1 $\uparrow$ A-SDF mu2021sdf 0.007 0.034 0.056 0.012 0.021 0.023 0.168 0.001 0.010 0.004 0.034
PARIS liu2023paris 0.221 0.151 0.319 0.091 0.423 0.024 0.421 0.530 0.031 0.032 0.224
D-3DGS yang2024deformable 0.064 0.180 0.125 0.073 0.168 0.195 0.233 0.131 0.022 0.023 0.121
ArtGS liu2025building 0.264 0.255 0.454 0.125 0.460 0.055 0.426 0.292 0.062 0.045 0.244
REArtGS (Ours) 0.254 0.215 0.577 0.186 0.465 0.052 0.509 0.608 0.065 0.050 0.298
EMD $\downarrow$ A-SDF mu2021sdf 2.123 0.951 0.955 1.606 1.684 1.878 1.731 1.115 1.620 1.837 1.559
PARIS liu2023paris 2.454 0.457 1.533 0.451 0.780 0.655 1.491 0.770 1.183 1.354 1.113
D-3DGS yang2024deformable 3.224 0.612 0.665 0.481 0.758 1.269 0.291 0.436 1.374 1.206 1.032
ArtGS liu2025building 2.207 0.562 0.412 0.563 0.525 1.171 0.631 0.724 1.297 0.750 0.884
REArtGS (Ours) 1.594 0.494 0.251 0.473 0.459 0.900 0.456 0.353 1.033 0.689 0.670
4 Experiments
4.1 Experimental Setting
Datasets. We conduct surface reconstruction experiments on synthetic dataset PartNet-Mobility xiang2020sapien and real-world dataset AKB-48 liu2022akb to evaluate the reconstruction quality of our method. Please see the Appendix for more details.
Metrics. Chamfer Distance (CD), F1-score and Earth Mover’s Distance (EMD) zhang2022deepemd are usually adopted as evaluation metrics for surface quality. However, these methods typically focus on the whole surface, endowing the unseen areas (such as the bottom of the object) with considerable weights during evaluation. This is unreasonable since we tend to be more concerned with the upper hemisphere regions of articulated objects. Therefore, we sample point clouds from the fragments between the rays cast by test cameras and the surface mesh to calculate EMD and CD, defined as CD (rs). We also evaluate the F1-score and CD for the whole surface, defined as CD (ws), to comprehensively evaluate the surface quality. Note that the CD values are multiplied by 1000 and the distance threshold of F1-score is set to 0.4. The calculation of these metrics can be referred to in our Appendix.
All the experiments are conducted on a single RTX 4090 GPU. We also provide detailed implementation of our approach and baselines in the Appendix.
4.2 Mesh Reconstruction Performance
We report the quantitative results of mesh reconstruction in Table. 1 and qualitative results are shown in Fig. 4. Note that we implement ArtGS without depth supervision for a fair comparison. As can be observed, our REArtGS outperforms state-of-the-art approaches in the mean across all metrics with 3.79, 1.236, 0.294 and 0.695 respectively. Compared to A-SDF mu2021sdf and Ditto jiang2022ditto that exploits 3D inputs for articulated object surface reconstruction, our method still achieves higher reconstruction quality in the majority of categories. In terms of PARIS, GOF and ArtGS, our approach exhibits significantly smoother and clearer surfaces as shown in Fig. 4, benefiting from our enhanced geometric constraints.
4.3 Mesh Generation Performance
We report the quantitative results of mesh generation in Table. 2 and present the qualitative results Fig. 5. Please refer to our Appendix for more qualitative results. Our REArtGS achieves the best mean results across all metrics with 4.10, 1.063, 0.298 and 0.670. Compared to PARIS, our method exhibits smoother surface generation results, as shown in Fig. 5. It is mainly because PARIS lacks a reasonable geometry initiation and directly utilizes the composite rendering, leading to the spatial overlap between the neural radiation fields of the start and end states. It can be also observed that our method is superior to ArtGS on most categories. This is primarily attributed to our high-quality geometry initialization from the reconstruction stage, as well as the motion-constrained deformable fields to yield dynamic generation.
Besides, we provide the part segmentation and joint parameter estimation results in the Appendix and  LABEL:sec:joint_parameter respectively.
Table 3: The ablation study of the unbiased SDF guidance on PartNet-Mobility dataset.
w/ SDF w/ Unbiased Reg. CD (ws) CD (rs) F1 EMD
5.96 2.242 0.250 1.366
4.38 1.921 0.273 0.817
3.79 1.236 0.294 0.695
Table 4: The ablation study of motion constraints on PartNet-Mobility dataset.
Settings CD (ws) CD (rs) F1 EMD
w/o motion constraints 18.65 2.270 0.209 1.105
w/ motion constraints 5.41 1.063 0.276 0.670
4.4 Ablation Studies
Table 5: The quantitative results on real-world AKB-48 dataset. We bold the best results and underline the second best results. * means we implement it without depth supervision for fair comparison.
Category Method Reconstruction Generation
[…]
Box PARIS liu2023paris 4.69 1.651 0.097 0.843 3.98 2.131 0.306 1.055
ArtGS liu2025building 8.61 0.967 0.143 0.697 10.23 1.606 0.309 0.851
REArtGS (Ours) 2.49 0.898 0.205 0.671 1.28 1.425 0.559 0.845
Stapler PARIS liu2023paris 0.23 0.104 0.551 0.165 0.47 0.198 0.540 0.194
ArtGS liu2025building 51.04 1.364 0.476 0.827 5.13 0.527 0.533 0.514
REArtGS (Ours) 0.24 0.037 0.728 0.137 0.34 0.044 0.582 0.151
Scissor PARIS liu2023paris 0.18 0.302 0.710 0.314 0.18 0.329 0.424 0.310
ArtGS liu2025building 11.43 0.073 0.588 0.192 5.53 0.089 0.600 0.213
REArtGS (Ours) 0.08 0.017 0.902 0.093 0.08 0.003 0.899 0.042
Cutter PARIS liu2023paris 2.56 1.371 0.570 0.302 45.02 1.684 0.309 1.183
ArtGS liu2025building 5.65 0.453 0.644 0.477 65.64 1.298 0.093 0.806
REArtGS (Ours) 0.13 0.016 0.874 0.090 51.56 1.559 0.228 0.884
Drawer PARIS liu2023paris 30.63 4.568 0.062 1.577 33.94 3.865 0.091 1.430
ArtGS liu2025building 19.41 7.478 0.086 1.935 12.60 1.624 0.094 0.911
REArtGS (Ours) 10.73 2.607 0.138 1.142 11.71 1.739 0.209 0.933
Eyeglasses PARIS liu2023paris 0.16 1.582 0.637 0.212 0.18 1.550 0.609 1.284
ArtGS liu2025building 27.49 1.384 0.557 0.832 12.12 1.592 0.747 0.955
REArtGS (Ours) 0.04 0.018 0.969 0.096 0.04 1.568 0.974 0.886
Mean PARIS liu2023paris 6.41 1.336 0.438 0.569 13.96 1.626 0.380 0.707
ArtGS liu2025building 20.61 1.953 0.416 0.826 18.54 1.123 0.396 0.708
REArtGS (Ours) 2.86 0.599 0.636 0.372 10.83 1.056 0.575 0.624
We conduct an ablation of the unbiased SDF guidance on PartNet-Mobility dataset to validate its effectiveness. The ablation setting and results are reported in Table. 3. The quantitative results prove that both SDF guidance and the unbiased regularization of SDF indeed improve surface reconstruction quality. This is primarily because the SDF representation establishes explicit geometric association between the opacity fields and scene surface through $\Phi_{k}$, and the unbiased SDF regularization further optimizes the distribution of Gaussian opacity fields.
We also conduct an ablation study of the motion constraints on PartNet-Mobility dataset and present the quantitative results in Table. 4. For the baseline without motion constraints, we employ three MLPs to learn the variations of spatial positions, scales, and quaternions respectively for Gaussian primitives, which is similar to D-3DGS. The significantly superior results demonstrate the effectiveness of the motion constraints.
Figure 6: The qualitative results of surface reconstruction on real-world AKB-48 liu2022akb repository. We show both textured and non-textured meshes for two-part and multi-part articulated objects.
4.5 Generalization to the Real World
We conduct reconstruction and generation experiments in the real world to investigate the generalization capacity of our REArtGS. We report the quantitative results comparing with PARIS and ArtGS in Table. 5, and present the qualitative results in Fig. 6. Our approach significantly outperforms PARIS and ArtGS in mean performance across all metrics. This demonstrates that our method exhibits strong generalization capability for real-world objects and enables high-quality surface reconstruction and generation as shown in Fig. 6.
5 Conclusion and Future Work
We propose REArtGS, a novel framework that achieves high-quality textured mesh reconstruction and dynamic generation of articulated objects using only RGB views of two arbitrary states. We propose an unbiased SDF guidance to regularize Gaussian opacity fields and model the Gaussian deformation fields constrained by kinematic structures for generating meshes of unseen states. Experiments show the superior performance of our method on both synthetic and real-world data.
The limitations of REArtGS lie on the requirement of camera pose prior and challenges of objects with transparent materials. Future works will introduce relative pose estimation to alleviate the dependence on camera poses and employ physically-based networks to model the transparent materials.
Acknowledgements
This work was supported by National Natural Science Foundation of China under Grant 62302143, 62576130, and Anhui Province’s Key Science and Technology Project 202423k09020007.